# A Generic Trajectory Planning Method for Constrained All-Wheel-Steering Robots

### v2

@@ Abstract
This paper presents a trajectory planning method for wheeled robots with fixed steering axes while the steering angle of each wheel is constrained. In the past, All-Wheel-Steering (AWS) robots, incorporating modes such as rotation-free translation maneuvers, in-situ rotational maneuvers, and proportional steering, exhibited inefficient performance due to time-consuming mode switches. This inefficiency arises from wheel rotation constraints and inter-wheel cooperation requirements. The direct application of a holonomic moving strategy can lead to significant slip angles or even structural failure. Additionally, the limited steering range of AWS wheeled robots exacerbates nonlinearity issues, thereby complicating control processes. To address these challenges, we developed a novel planning method termed Constrained AWS (C-AWS), which integrates second-order discrete search with predictive control techniques. Experimental results demonstrate that our method adeptly generates feasible and smooth trajectories for C-AWS while adhering to steering angle constraints. Code and video can be found at https://github.com/Rex-sys-hk/AWSPlanning.
@@ I-A Motivation
The benefits of the omnidirectional wheel chassis, particularly during low-speed maneuvers, are well-recognized, enabling operations like spot turning and translation that are unfeasible with traditional platforms with Ackerman geometry [1, 2]. However, the intricacy and lack of intuitiveness in the human-machine interaction for All-Wheel-Steering (AWS) vehicles have hampered their widespread usage [3].
@@
Figure 1: (a) Depicts a typical autonomous airport baggage carrier, equipped with LiDAR, cameras, computing units and AWS chassis. (b) Illustrates the design of the vehicle’s steering wheel. The image reveals that despite the wheel’s ability to rotate via the top-mounted steering motor, its free rotation is constrained by essential circuits and hydraulic systems. (c) and (d) demonstrate the operation of the AWS vehicle through mode switching and the movement optimization achieved by the proposed C-AWS algorithm, respectively.
@@
The advent of autonomous driving technologies is improving the functionality of all-wheel-steering robots, making their control more manageable. Autonomous operation requires consideration of the rotation angles and speeds of all wheels, introducing complexities absent in the simplified motion control of Ackerman geometry [4]. The concept of the Instantaneous Center of Motion (ICM) [5], is widely employed in describing the motion of rigid bodies, to make sure coherence in the motion states of each wheel. ICM theory posits that the motion of a rigid body can be simplified as rotational motion around a specific center, characterized by angular velocity and radius. However, the ICM model exhibits two singular conditions when rotating radius approaching 0 or infinity, potentially leading to scenarios where the controller lacks a viable solution or may even crash. Christian et al. [6] remapped the rotation radius to a periodic function using spherical coordinates to address the singularity problem of the ICM method when the rotating radius approaches infinity. Nevertheless, Christian et al. only solves the mapping relationship between moving speed, wheel walking speed, and steering angle, but cannot solve the robot motion control problem. This method does not ensure that the vehicle can effectively follow a path. Reza et al. [7, 8] comprehensively defined the path-following problem for AWS robots, considering wheel velocity constraints. They estimated the maximum speed for each wheel to achieve time optimality and considered the remaining trajectory length to prioritize either distance minimization or angle error reduction. However, their model does not address acceleration constraints, leading to frequent wheel oscillations and significant acceleration when tracking specific directional trajectories. In addition, AWS platforms have problems with steering constraints. For example, airport baggage carriers confined to a steering range of ±$\pm 90$° [9], attributed to factors like structural durability, hydraulic and electric system requirements, as illustrated in fig. 1(a),(b). Traditionally, to circumvent constraints, practitioners treated motion within a specific constraint range as a distinct driving mode and manually switched between these modes, leading to a reduction in movement flexibility for the AWS chassis as demonstrated in fig. 1(c).
@@ I-B Contribution
To enhance the operating efficiency of Constrained AWS(C-AWS) platforms, smoothness, and reliability under constrained conditions, we adopt a predictive control strategy. Our approach involves initially identifying a viable trajectory to the target pose and subsequently refining it to minimize travel time within given constraints. The trajectory search accounts for both the feasibility and cost related to wheel movement, as well as the connectivity within the configuration space. Consequently, we introduced the time-cost-2nd-order hybrid A* [10] method for initial trajectory searching, which evaluates wheel conditions and integrates traveling cost into time consumption. In this way, the planner can search for efficient driving mode priors instead of manually defining them. In the optimization process, our method referred to the Model Predictive Control (MPC) framework, emphasizing the need for a motion state equation that is third-order differentiable and continuous to adhere to acceleration constraints. Given that the control model by Christian et al. [6] exhibits singularity in steering angle and lacks differentiability when rotation radius is 0, we integrated a differentiable rigid body motion description [11] with steering constraints. This integration was facilitated by introducing intermediary variables to circumvent the robot state singularities problem and linearizing the constraint expressions.
@@
    The time-cost-2nd-order initial trajectory planner, which is widely suitable for robots with fixed steering axis positions.
@@
    The universal smoother that incorporates our kinematic model, linearizing steering constraints to generate feasible paths and control mechanisms for C-AWS.
@@
    Digital twinning of the experimental platform and the conduction of experiments in a simulation environment verifying the effectiveness of our method.
@@ II-A Kinematic Models
Kinematic models are represented as ordinary differential equations. These equations describe the relationship between the changing states and their rates of change over time. Mass point models are a simplified representation where moving states are linearly additive, like when the heading of the controlled object is independent of its motion. This model is suitable for holonomic moving platforms equipped with Macnumm wheels [12]. Control problems of such models can be classified as Linear Programming (LP) problems and can usually be solved with simple negative feedback control or Linear Quadratic Regulator (LQR). Nonlinear dynamic models are used to describe physical systems with dependent states. Controlling with this kind of model typically involves numerical Nonlinear Programming (NLP) solving methods. For instance, single-wheel models are used on differential steering wheeled robots [13], which adjust the speed difference between drive wheels to change the facing direction of the vehicle and adjust its moving direction. The bicycle model is another nonlinear model simplifying the mathematical representation of the dynamics of vehicles with Ackerman geometry. This model captures the essential motion features of vehicles like bicycles, whose rear wheels cannot be steered or whose rear wheels are steered in a certain proportion [14], and even trailers [15]. The steering angle has an injective relationship to the curvature of an arc tangent to the heading direction at a specific point. In contrast to the aforementioned nonlinear dynamic models, AWS robots allow for dynamic adjustment of the rotating center. This transforms the control of lateral movement from one-dimensional to two-dimensional, significantly enhancing the robot’s mobility. However, in the planning and control of AWS robots, the relationship between the wheel’s steering angle and the motion state is non-injective and is further complicated by substantial inter-wheel constraints. These issues amplify the non-linearity and discontinuity of the model, rendering it unsuitable for direct resolution through NLP solvers. In our C-AWS algorithm, we reformulated this problem to enable its solution using Nonlinear Programming (NLP) solvers.
@@ II-B Trajectory Planning
Trajectory planning is a process that generates a feasible and followable path for a robot, assigning specific speed and pose requirements at certain time steps. This process is governed by two core criteria: ensuring the trajectory is collision-free and trackable [16]. Sampling-based methods are popular in robotics for finding collision-free paths due to their versatility in incorporating user-defined objectives. These methods encompass state lattice [17, 18] approaches and probabilistic planners. They essentially operate by sampling the robot’s possible states within the configuration space to discover a viable trajectory between the start and goal nodes. Lattice-based planners break down the continuous state space into a discrete lattice graph structure for planning purposes. Subsequently, algorithms such as Dijkstra’s algorithm are applied to identify the optimal trajectory through this graph. Excluding lattice planners, probabilistic planners exemplified by Rapidly-exploring Random Trees (RRT) [19] generate a feasible path by iteratively growing a tree of states starting from the initial node. The sampling process inherent in these methods [20, 21, 10] accounts for the vehicle’s dynamic constraints. In the process of searching for paths, piecewise parametric curves are extensively used. For example, Dubins’ paths [22] and Reed-Shepp’s paths [23] decompose complex paths into sequences of simpler maneuvers, such as constant turning and straight-line motion. Meanwhile, spiral curves facilitate gradual changes in direction and speed [24]. Specifically, the forward movement of bicycles can be modeled as an arc, ensuring feasibility for non-holonomic models. Therefore, our time-cost-2nd-order trajectory searching method has been developed based on the hybrid A* algorithm which can efficiently take advantage of this feature.
MPC techniques address trajectory planning by conceptualizing it as an Optimal Control Problem (OCP), subsequently transformed into a Nonlinear Programming (NLP) problem [25, 26, 27]. Regarding the functionality of MPC, our algorithm employs this framework with meticulous model design and high-quality initial guesses. Trajectory optimization for vehicles at each timestep utilizes differentiable curves or numerical methods to maintain state continuity. Differentiable curves like splines are generated by interpolating between control points [28] while polynomials are determined by coefficients [29]. However, the use of parameterized curves introduces bias related to curve property or degree of freedom. To solve these trajectory optimization problems numerically, methods such as Euler [30] and advanced Runge-Kutta (RK) algorithms are employed in our method. Specifically, we utilize the RK4 algorithm, which strikes a balance between accuracy and computational efficiency.
@@ III-A Rigid Body Vehicle Model
The location of a rigid body in 2D space can be represented by SE(2)
$\begin{bmatrix}\mathbf{R}&\mathbf{p}\\ 0&1\end{bmatrix}\quad\text{where}\quad\mathbf{R}\in SO(2)\quad\text{and}\quad\mathbf{p}\in\mathbb{R}^{2},$
where the rotation matrix and position vector are
$\mathbf{R}(\theta)=\begin{bmatrix}\cos{\theta}&-\sin{\theta}\\ \sin{\theta}&\cos{\theta}\end{bmatrix},\mathbf{p}=\begin{bmatrix}x\\ y\end{bmatrix}.$
@@
Figure 2: The schematic representation of an AWS robot kinematic model. Please note the $\omega$ direction only indicates the positive direction, not the rotating direction at the instance.
@@
and time derivative of $\delta_{w}$ is,
$\begin{split}\frac{\partial\bm{\delta}_{w}}{\partial t}&=\frac{\partial\arctan_{2}{(v_{wy},v_{wx})}}{\partial t}\\ &=\frac{-v_{wy}\dot{v}_{wx}v_{w}+v_{wx}\dot{v}_{wy}v_{w}-v_{wx}v_{wy}-v_{wy}^{2}+v_{w}^{2}\dot{v}_{wy}}{2v_{w}^{2}(v_{w}+v_{wx})}.\end{split}$ (6)
Nevertheless, it is not used in the optimizing process together with eq. 3, eq. 4 to avoid discontinuity and singularity.
@@ III-B Path Initialization
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 4: The sampled maneuvers with $\delta_{lim}=$ 90°,75°,60°and front-75°-rear-0°respectively. The red points are the end of the velocity vector, the blue points are the end of trajectories, and the black points are instantaneous centers. In all the forward samples above, N of $\mathcal{E}$ is 8, N of $\Psi$ is 8, and $\omega$ are uniformly sampled 8 times in range $[-\pi/2,\pi/2]$.
To offer a high-quality feasible route for the optimizer, a global planning method based on HybridA* is developed. In our design, not only the first-order states like position and heading are sampled, but also the body velocity is sampled and recorded. We simplify the integration process from high-order variables to low-order variables, which is a tradeoff in calculation time and accuracy. For the forward simulation process, we customized the grid sampling method for this all-wheel steering chassis. To encourage that each sampling explores new states, we conduct grid sampling of ICM on the spherical coordinate. The feasible ICM location regarding steering limits is shown in Fig. 3. It must be within the intersection of the areas that the normals of the movement directions of all wheels can pass through.
@@
$\begin{split}\mathbf{r}&=\begin{bmatrix}-\tan{\varepsilon}\cdot\cos{\psi}\\ -\tan{\varepsilon}\cdot\sin{\psi}\end{bmatrix},\forall\varepsilon\in\mathcal{E},\forall\psi\in\Psi,\\ \text{subject to},\\ &\underaccent{\bar}{\delta}_{w}\leq|\arctan_{2}(\mathbf{r}_{y,ICM,w},\mathbf{r}_{x,ICM,w})|-\frac{\pi}{2}\leq\bar{\delta}_{w}\\ &\mathbf{r}_{ICM,w},\bar{\delta}_{w}\in\mathcal{D}_{ulim},\underaccent{\bar}{\delta}_{w}\in\mathcal{D}_{llim},\forall\mathbf{w}\in\mathbf{W}.\\ \end{split}$
The rotation speed $\omega$ can be directly sampled in a certain range. However, as $r$ increases, certain rotation speeds may cause the traveling distance to be too long, so we set the longest arc length to control the farthest traveling distance of each step. The sampled $\omega$ can be confined by a maximum curvature length. The sampling results of different steering limits are demonstrated in fig. 4. Through simple mathematical derivation according to eq. 1, we can obtain the current speed $\Delta v$ and next position $\Delta x$ after a certain time under this r and omega. Also, we can derive the relative wheel speed $v_{w,n}$ and steering angle $\delta{w,n}$ of that node according to eq. 2, and eq. 4. It is worth noting that for the common situation where the limit angle is less than 90 degrees, there is a single mapping relationship between the wheel angle speed and the wheel friction point speed. If it exceeds 90 degrees, an additional mapping table will be needed. We will not discuss this for the time being. We add the speed and position values as attributes of the next point to the new node constructed during the hybrid A* search. In order to better utilize the speed attribute of the node, we express the cost in time to unify the cost function and avoid weight adjustment of speed and distance information. So the traveling cost is computed by,
@@
where $k_{h}$ is the hypothesis coefficient, which represents the ratio of the time consumption of locomotion to the time consumption of the vehicle’s motion with the maneuver. The algorithm is the same as ordinary hybrid A* in the unmentioned aspects. The discretization search process of initial guess will also output the wheel steering and vehicle discontinuous state space flags, which will be explained in the next section.

### v3

@@ Abstract
This paper presents a generic trajectory planning method for wheeled robots with fixed steering axes while the steering angle of each wheel is constrained. In the existing literatures, All-Wheel-Steering (AWS) robots, incorporating modes such as rotation-free translation maneuvers, in-situ rotational maneuvers, and proportional steering, exhibit inefficient performance due to time-consuming mode switches. This inefficiency arises from wheel rotation constraints and inter-wheel cooperation requirements. The direct application of a holonomic moving strategy can lead to significant slip angles or even structural failure. Additionally, the limited steering range of AWS wheeled robots exacerbates non-linearity characteristics, thereby complicating control processes. To address these challenges, we developed a novel planning method termed Constrained AWS (C-AWS), which integrates second-order discrete search with predictive control techniques. Experimental results demonstrate that our method adeptly generates feasible and smooth trajectories for C-AWS while adhering to steering angle constraints. Code and video can be found at https://github.com/Rex-sys-hk/AWSPlanning.
@@ I-A Motivation
The benefits of the omnidirectional wheel chassis, particularly during low-speed maneuvers, are well recognized, enabling operations such as spot turning and translation that are unfeasible with traditional platforms with Ackerman geometry [1, 2]. However, complexity and lack of intuitiveness in human-machine interaction for All-Wheel-Steering (AWS) vehicles have hampered their widespread use [3].
@@
Figure 1: (a) A typical autonomous airport baggage carrier, equipped with LiDAR, cameras, computing units, and AWS chassis. (b) Design of the vehicle’s steering wheel. The image reveals that despite the wheel’s ability to rotate via the top-mounted steering motor, its free rotation is constrained by essential circuits and hydraulic systems. (c) and (d) The operation of the AWS vehicle through mode switching and the movement optimization achieved by the proposed C-AWS algorithm, respectively.
@@
The advent of autonomous driving technologies is improving the functionality of all-wheel-steering robots, making their control more manageable. Autonomous operation requires consideration of the rotation angles and speeds of all wheels, introducing complexities absent in the simplified motion control of Ackerman geometry [4]. The concept of the Instantaneous Center of Motion (ICM) [5], is widely employed in describing the motion of rigid bodies, to make sure coherence in the motion states of each wheel. ICM theory posits that the motion of a rigid body can be simplified as rotational motion around a specific center, characterized by angular velocity and radius. However, the ICM model exhibits two singular conditions when rotating radius approaching zero or infinity, potentially leading to scenarios where the controller lacks a viable solution or may even crash. Previous works [6, 7] avoid singularities by spliting the action space into modes which would prevent the robot from making full use of its flexibility. Christian et al. [8] remapped the rotation radius to a periodic function using spherical coordinates to address the singularity problem of the ICM method when the rotating radius approaches infinity. Nevertheless, Christian et al. only solves the mapping relationship between moving speed, wheel walking speed, and steering angle, but cannot solve the robot motion control problem. This method does not ensure that the vehicle can effectively follow a path. Reza et al. [9, 10] comprehensively defined the path-following problem for AWS robots, considering wheel velocity constraints. They estimated the maximum speed for each wheel to achieve time optimality and considered the remaining trajectory length to prioritize either distance minimization or angle error reduction. However, their model does not address acceleration constraints, leading to frequent wheel oscillations and significant acceleration when tracking specific directional trajectories. In addition, AWS platforms have problems with steering constraints. For example, airport baggage carriers are confined to a steering range of ±$\pm 90$° [11], attributed to factors like structural durability, hydraulic and electric system requirements, as illustrated in Fig. 1(a),(b). Traditionally, to circumvent constraints, practitioners treated motion within a specific constraint range as a distinct driving mode and manually switched between these modes, leading to a reduction in movement flexibility for the AWS chassis as demonstrated in Fig. 1(c).
@@ I-B Contribution
To enhance the operating efficiency of Constrained-AWS (C-AWS) platforms, smoothness, and reliability under constrained conditions, we adopt a predictive control strategy. Our approach involves first identifying a viable trajectory to the target pose and subsequently refining it to minimize travel time within given constraints. The trajectory search accounts for both the feasibility and the cost related to wheel movement, as well as connectivity within the configuration space. Consequently, we introduced the time-cost-2nd-order Hybrid-A* [12] method for initial trajectory search, which evaluates wheel conditions and integrates travel cost into time consumption. In this way, the planner can search for efficient priors of the driving mode instead of manually defining them. In the optimization process, our method referred to the Model Predictive Control (MPC) framework, emphasizing the need for a motion state equation that is differentiable by third order and continuous to adhere to the acceleration constraints . Given that the control model by Christian et al. [8] exhibits singularity in steering angle and lacks differentiability when rotation radius is zero, we integrated a differentiable description of rigid body motion [13] with steering constraints. This integration was facilitated by introducing intermediary variables to circumvent the robot state singularities problem and linearizing the constraint expressions.
@@
    We present a time-cost-2nd-order initial trajectory planner, which is widely suitable for robots with fixed steering axis positions.
@@
    We propose a new universal smoother that incorporates our kinematic model, linearizing steering constraints to generate feasible paths and control mechanisms for C-AWS.
@@
    We demonstrate the digital twinning of the experimental platform and the conduction of experiments in a simulation environment verifying the effectiveness of our method.
@@ II-A Kinematic Models
Kinematic models are represented as ordinary differential equations. These equations describe the relationship between the changing states and their rates of change over time. Mass point models are a simplified representation where moving states are linearly additive. Control problems of such models can be classified as Linear Programming (LP) problems and can usually be solved with simple negative feedback control or minimizing linear quadratic cost functions. Non-holonomic dynamic models are used to describe physical systems with dependent states. Controlling with this kind of model typically involves numerical Nonlinear Programming (NLP) solving methods [14, 15]. For instance, single-wheel models are used on differential steering wheeled robots [16], which adjust the speed difference between drive wheels to change the facing direction of the vehicle and adjust its moving direction. The bicycle model is another non-linear model that simplifies the mathematical representation of the dynamics of vehicles with Ackerman geometry. This model captures the essential motion features of vehicles such as bicycles, whose rear wheels cannot be steered or whose rear wheels are steered in a certain proportion [17]. The steering angle has an injective relationship to the curvature of an arc tangent to the heading direction at a specific point. In contrast to the aforementioned nonlinear dynamic models, AWS robots allow for dynamic adjustment of the rotating center. This transforms the control of lateral movement from one-dimensional to two-dimensional, significantly enhancing the robot’s mobility. However, in the planning and control of AWS robots, the relationship between the wheel’s steering angle and the motion state is non-injective and is further complicated by substantial inter-wheel constraints. These issues amplify the non-linearity and discontinuity of the model, rendering it unsuitable for direct resolution through NLP solvers. In our C-AWS algorithm, we reformulated this problem to enable its solution using NLP solvers.
@@ II-B Trajectory Planning
Trajectory planning is a process that generates a feasible and followable path for a robot, assigning specific speed and pose requirements at certain time steps. This process is governed by two core criteria: ensuring that the trajectory is collision-free and trackable [18]. Sampling-based methods are popular in robotics for finding collision-free paths due to their versatility in incorporating user-defined objectives. These methods encompass state lattice [19, 20] approaches and probabilistic planners. They essentially operate by sampling the robot’s possible states within the configuration space to discover a viable trajectory between the start and goal nodes. Lattice-based planners break down the continuous state space into a discrete lattice graph structure for planning purposes. Subsequently, algorithms such as Dijkstra’s algorithm are applied to identify the optimal trajectory through this graph. Excluding lattice planners, probabilistic planners exemplified by Rapidly-exploring Random Trees (RRT) [21] generate a feasible path by iteratively growing a tree of states starting from the initial node. The sampling process inherent in these methods [22, 23, 12] accounts for the dynamic constraints of the vehicle. In the process of searching for paths, piecewise-parametric curves are extensively used. Specifically, the forward movement of bicycles can be modeled as an arc, ensuring feasibility for the non-holonomic model in Hybrid-A*[12]. Therefore, our time-cost-2nd-order trajectory searching method has been developed based on the Hybrid-A* algorithm, which can efficiently take advantage of this feature.
MPC techniques address trajectory planning by conceptualizing it as an Optimal Control Problem (OCP), subsequently transformed into an NLP problem [24, 25, 26]. Regarding the functionality of MPC, our algorithm employs this framework with meticulous model design and high-quality initial guesses. Trajectory optimization for vehicles at each time step utilizes differentiable curves or numerical methods to maintain state continuity. Differentiable curves, such as splines, are generated by interpolating between control points [27], while polynomials are determined by coefficients [28]. However, the use of parameterized curves introduces bias related to the curve property or degree of freedom. To solve these trajectory optimization problems numerically, we employ methods such as Euler [29] and advanced Runge-Kutta (RK) algorithms. Specifically, we utilize the RK4 algorithm, which strikes a balance between accuracy and computational efficiency.
@@ III-A Rigid Body Vehicle Model
The pose $\mathcal{P}$ of a rigid body in 2D space can be represented by 2 dimentional Special Euclidean Group $SE(2)$, i.e.,
$\mathcal{P}=\begin{bmatrix}\mathbf{R}&\mathbf{p}\\ 0&1\end{bmatrix},$
where the rotation $\mathbf{R}$ is belong to 2 dimentional Special Othogonal Group $SO(2)$ and position $\mathbf{p}$ is a 2 dimentional vector, i.e.,
$\mathbf{R}(\theta)=\begin{bmatrix}\cos{\theta}&-\sin{\theta}\\ \sin{\theta}&\cos{\theta}\end{bmatrix}\in SO(2)\text{,~{}}\mathbf{p}=\begin{bmatrix}x\\ y\end{bmatrix}\in\mathbb{R}^{2}.$
@@
Figure 2: The schematic representation of an AWS robot kinematic model. Note that the $\omega$ direction only indicates the positive direction, instead of the rotating direction at the instance.
@@ III-B Path Initialization
Figure 4: The sampled maneuvers with $\delta_{lim}=$ 90°, 75°, 60° and front-75°-rear-0° for (a), (b), (c) and (d) respectively. The red points are the end of the velocity vector, the blue points are the end of trajectories, and the black points are instantaneous centers. In all the forward samples above, $N$ of $\mathcal{E}$ is 8, $N$ of $\Psi$ is 8, and $\omega$ is uniformly sampled 8 times in range $[-\pi/2,\pi/2]$.
To offer a high-quality feasible route for the optimizer, a global planning method based on Hybrid-A* is developed. This will reduce the failure of the NLP solver and increase the solving speed. In our design, not only the first-order states like position and heading are sampled, but also the body velocity is sampled and recorded. We simplify the integration process from high-order variables to low-order variables, which is a trade-off in calculation time and accuracy. For the forward simulation process, we customized the grid sampling method for this all-wheel steering chassis. To encourage that each sampling explores new states, we conduct grid sampling of ICM on the spherical coordinate to deal with singularity when steering is zero and align with the sequential optimization problem. The feasible ICM location regarding steering limits is shown in Fig. 3. It must be within the intersection of the areas that the normals of the movement directions of all wheels can pass through.
@@
$\begin{split}\mathbf{r}&=\begin{bmatrix}-\tan{\varepsilon}\cdot\cos{\psi}\\ -\tan{\varepsilon}\cdot\sin{\psi}\end{bmatrix},\forall\varepsilon\in\mathcal{E},\forall\psi\in\Psi,\\ \text{subject to},\\ &\underaccent{\bar}{\delta}_{w}\leq|\arctan_{2}(\mathbf{r}_{y,ICM,w},\mathbf{r}_{x,ICM,w})|-\frac{\pi}{2}\leq\bar{\delta}_{w}\\ &\mathbf{r}_{ICM,w},\bar{\delta}_{w}\in\mathcal{D}_{ulim},\underaccent{\bar}{\delta}_{w}\in\mathcal{D}_{llim},\forall w\in\mathbf{W}.\\ \end{split}$
The rotation speed $\omega$ can be directly sampled in a certain range. However, as $r$ increases, certain rotation angular regarding that time step may cause the traveling distance to be too long, so we set the longest curvature length to control the farthest traveling distance of each time step. The sampling results of different steering limits are demonstrated in Fig. 4. Through simple mathematical derivation according to Eq. (1), we can obtain the current speed $\Delta v$ and next position $\Delta x$ after a certain time under this r and omega. Also, we can derive the relative wheel speed $v_{w,n}$ and steering angle $\delta{w,n}$ of that node according to Eq. (2), and Eq. (4). It is worth noting that for the common situation where the limit angle is less than 90°, there is a single mapping relationship between the wheel angle speed and the wheel friction point speed. If it exceeds 90°, an additional mapping table will be needed. We will not discuss this for the time being. We add the speed and position values as attributes of the next point to the new node constructed during the Hybrid-A* search. In order to better utilize the speed attribute of the node, we express the cost in time to unify the cost function and avoid weight adjustment of speed and distance information. So the traveling cost is computed by,
@@
where $k_{h}$ is the hypothesis coefficient, which represents the ratio of the time consumption of locomotion to the time consumption of the vehicle’s motion with the maneuver. The algorithm is the same as ordinary Hybrid-A* in the unmentioned aspects. The discretization search process of initial guess will also output the wheel steering and vehicle discontinuous state space flags, which will be explained in the next section.